# Stochastic Adaptive Estimation in Polynomial Curvature Shape State Space for Continuum Robots

### v4

@@ Abstract
Index Terms:
Shape estimation, Continuum robots, Extended Kalman filter, Interacting multiple model, polynomial curvature kinematics, Observability analysis
@@ I Introduction
Inspired by biological features such as an elephant’s trunk or an octopus’s tentacles, continuum robots are an unconventional class of robots that have inherent large degrees of flexibility, underactuation, and compliance. Such "soft" features allow them to have safe physical interactions with the environments[1]. With recent advances in instrument miniaturization, continuum robot designs have been widely used in many minimally invasive surgery applications [2, 3], and their advantages are further leveraged for safe manipulation and access to deep anatomy.
@@ I-A Sensing Technique
Shape sensing is pivotal in this context by supplying the observational data critical for robust shape estimation. A diverse array of sensing techniques is deployed to acquire the shape information. External sensing involves equipping robots with sensors such as Electromagnetic (EM) sensors[4, 5, 6, 7, 8, 9], vision-based systems[10, 7, 11], or Inertial Measurement Units (IMUs)[12, 13] to obtain pose information externally. Alternatively, intrinsic sensing leverages internal sensors to enhance joint-level information, utilizing encoders or passive cable displacement measurements [14]. Given their infinite degrees of freedom (DoF), the challenge with continuum robots lies in capturing comprehensive pose information along their entire length with a finite number of sensors. While technologies like Fiber Bragg Grating (FBG) sensors[15, 16, 17, 18, 19, 9] and vision-based systems can map out the full robot shape, constraints such as budget and the specific requirements of the application often preclude their ubiquitous use.
@@ I-B Representations in Modeling
Different shape representation methods yield various formulations of the backbone curvature, which directly impact continuum robots’ computed position and orientation. Representation methods can be categorized as distributed and lumped parameterizations for the backbone curvature[20, 21].
Distributed parameterization methods represent the backbone curvature as a continuous function of the arc length, characterized by an infinite number of parameters. These methods assume that the curvature distribution is variable along the length of the robot. A classic example is the Cosserat rod theory (CRT), which employs distributed parameter modeling to account for the underlying mechanics by solving for positional, force, and moment boundary conditions [22, 23, 24, 25, 26].
Lumped parameterization uses a finite set of parameters to describe the backbone shape of continuum robots geometrically. This approach primarily includes three methods: the Pseudo-Rigid-Body Model (PRBM), the constant curvature (CC) assumption, and the modal approach[20]. PRBM approximates the central backbone shape of continuum robots as a series of rigid links connected by torsion springs[27, 28, 29, 30]. The CC model assumes that each segment of the continuum robot follows a constant curvature, typically represented as a circular arc, with a single curvature parameter defining the shape. Due to its simplification and computation efficiency, it is deployed quickly for ideal situations such as uniform mass/material distribution, frictionless transmission, and free from external loading. As a widely adopted kinematics-based approach, The CC model is extensively used in applications ranging from surgical procedures [31, 32, 33, 34] and field operations[35] to underwater exploration[36] and aerial maneuvers [37]. In contrast, the modal approach[38] represents backbone curvature as a linear combination of modal shape functions (e.g., monomials[39], Euler curves [40], Chebyshev polynomials[14], and polynomial curvature kinematics (PCK)[12, 41]). A modal approach offers the advantage of capturing complex, continuous deformations with a relatively small number of parameters, making it computationally efficient while still maintaining high accuracy in shape representation.
In this work we adopt the polynomial-curvature (PCK) modal family as the lumped parameterization; our experiments bring prior PCK theory into hardware validation under limited/sparse sensing, responding to calls for experimental studies and fair comparisons among polynomial-shape methods [41, 42].
@@ I-C Estimation Approach
Various sensing techniques and representation methods drive multiple estimation approaches tailored to the specific scenario. In static configurations, where the continuum robot remains stationary and its state does not change over time, researchers often employ curve fitting or optimization methods based on modal approaches[15, 4, 40, 14], and PRBM[10]. Considering model uncertainty and measurement noise, Lilge et al. [6] and Ferguson et al. [43] implemented Gaussian Process Regression (GPR) for static shape estimation under arclength domain using CRT.
Dynamic configurations involve continuous state updates as the continuum robot moves and its state changes over time. In such scenarios, stochastic observer-based estimation methods, particularly filtering approaches, are preferred for their ability to optimize state estimation by balancing the influences of process and measurement models, where noise from both sources significantly impacts accuracy and robustness. Although the Extended Kalman Filter (EKF) is a general estimation technique that can be paired with various forward models, most existing continuum robot shape-estimation studies have employed EKF in conjunction with CC or piecewise constant curvature (PCC) models[8, 11, 13]. In addition to the EKF, alternative filtering methods—such as the Unscented Kalman Filter (UKF) and particle filters—have also been explored [44, 7]. Each technique presents trade-offs in computational complexity, robustness, and ease of implementation. For instance, UKF’s sigma-point method can improve estimation in highly nonlinear systems, whereas particle filters can capture non-Gaussian noise at the expense of increased computation time. In certain cases, the EKF remains a better choice when the system model is sufficiently accurate, or when computational resources are limited. Another well-established multi-model estimation approach is the Interacting Multiple Model (IMM) approach [45, 46]. Initially developed for tracking targets with Markovian switching dynamics, IMM simultaneously executes multiple Kalman-type filters (e.g., EKF or UKF), each associated with distinct motion models. Estimates from these parallel filters are integrated based on model probabilities updated via measurement innovations, commonly determined through likelihood or Normalized Innovation Squared (NIS) tests. IMM-based approaches are advantageous when a system switches between distinct kinematic regimes or experiences abrupt behavioral changes, conditions frequently encountered in continuum robotics due to varying loads, external interactions, or segmental bending transitions. Despite extensive IMM applications in mobile robotics and trajectory tracking, its deployment in dynamic shape estimation for continuum robots remains scarce, highlighting a valuable research direction in robust multi-model estimation.
@@ III-A Polynomial Curvature Kinematics
A precise representation and model of the shape for continuum robots([41],[12]) is rooted in PCK. Leveraging this model, all the poses spanning the arclength can be ascertained by integrating a curvature function. The general curvature function is described as an infinite expansion of monomials:
@@ III-A2 Forward kinematics of configuration-to-task space
In this paper, we follow the frame notation of [39] to derive the kinematics (illustrated in Fig. 2), and we list the definitions of all frames in Table I. Because we assume no torsional twist along the backbone, the robot’s shape can be fully described in a single plane (see the right side of Fig. 2). Accordingly, only one bending angle $\delta(t)$ is required to capture the in-plane motion, with Frame $\{1\}$ having its axes $\hat{\mathbf{x}}_{1}$ and $\hat{\mathbf{z}}_{1}$ aligned with the bending plane. Unless stated otherwise, all vectors and parameters are expressed in Frame $\{1\}$.
@@ III-B Shape State Representation
However, this model assumes uniform curvature along the robot’s length and simplifies the representation, which often does not accurately reflect the complex shape dynamics in environments with disturbances. In many practical deployments (e.g., endoluminal procedures or constrained workspaces), deformation is predominantly in-plane, and in-plane shape sensing is therefore critical for navigation and control. Contemporary systems and surveys report continued use of CC/PCC pipelines in applied settings [47, 48, 49, 50], while hardware strategies that raise torsional stiffness (thereby suppressing twist in normal use) further support effectively planar operation [51, 52]. Moreover, in multi-segment tendon-driven continuum robots the bending plane angle $\delta$ can be directly computed from cable displacements using established kinematic models [33] and thus does not need to be estimated as an additional state variable. In this work, since we focus exclusively on in-plane bending motions, we omit the bending plane angle $\delta$ from our analysis, thereby simplifying the model to better suit our specific research needs. We intentionally exclude the plane angle $\delta$ from our shape state to maintain a clear focus on how polynomial curvature coefficients map directly to sensor measurements. Incorporating $\delta$ requires augmenting every matrix and Jacobian with rotation operators, significantly increasing the notational overhead. This extra layer, though standard, can obscure the core novelty of using non-constant polynomial curvature within a stochastic observer. By omitting $\delta$, we simplify our presentation of the observer’s theoretical development while still enabling 3D motion in real applications by commanding $\delta$ externally, if desired.
@@ III-C State Transition Model
The scope of our work is confined to evaluating our observer-based shape estimation framework. Consequently, we do not incorporate advanced control strategies but instead opt for a straightforward open-loop control. For simplicity—and without loss of generality—we designate the control input $\mathbf{u}$ as the change in the bending angle at the tip, denoted by $\delta\theta_{e}$. This choice is supported by previous studies, such as Camarillo et al.[33], which demonstrate that in tendon-driven continuum manipulators, the tip bending angle is closely linked to the actuator commands. Similarly, Peng et al.[53] successfully used the bending angle as a control parameter in an aerial continuum manipulator, illustrating that this input is not only practical but also adaptable across different application domains. Moreover, our IMM-EKF framework is inherently flexible: once a reliable mapping between the chosen control input and the shape state (modal coefficients) is established, the estimation process can accommodate alternative control inputs if desired.
@@ IV-B1 Mitigating Bias in State Dimensionality
In the mixing stage of the IMM, the mixed initial state for each model is derived based on Equation (36), which requires uniform dimensionality across all state vectors and their corresponding covariance matrices. The conventional IMM approach often addresses dimensional discrepancies by setting augmented components in state vectors with smaller dimensions to zero. This leads to potential biases in the mixed initial states of models with larger dimensions. Inspired by [54], which provides an unbiased mixing solution for two models, we propose an unbiased mixing strategy tailored for multiple models with varying dimensions, illustrated using the case of $\mathbb{M}=\left\{M^{\mathit{(0)}},M^{\mathit{(1)}},M^{(2)}\right\}$.
@@ IV-B2 Designing Adaptive Transition Strategies
The conventional IMM method often struggles to meet performance expectations in dynamic environments due to its reliance on a fixed TPM predicated on prior knowledge. This limitation has spurred researchers to develop methods that enable the TPM to adapt more effectively, not just based on historical data but also incorporating real-time information. We build upon these efforts in our work, especially drawing on the correction functions proposed in [55] and [56]. We leverage these concepts to refine and implement a TPM adjustment mechanism that responds dynamically to current system states, enhancing the IMM’s responsiveness and accuracy in diverse operation scenarios.
@@
We utilize an activating correction function described in [56] to reduce lag errors and improve the responsiveness of IMM when the model jumps. The activating function is designed as
@@
Unlike the strategy of using $g_{k}^{(i)}$ at all time steps in [56], we introduce the minimum likelihood ratio between models as a prior judging condition to know if the current model jumps. The reason is that this correction function may cause instability in cases with large noises and no dominant model. The likelihood ratio is expressed as
@@ V Observability Analysis
For each model in our IMM framework, an EKF is used to estimate the shape state $\mathbf{m}$. Convergence of the EKF requires that the system be observable. In linear time-invariant (LTI) systems, observability is often checked via the rank of the observability matrix. In nonlinear systems, however, we adopt local weak observability [57], which relies on Lie derivatives to capture how variations in the shape state map to sensor outputs.
Mahoney et al. [58] also employ a rank-based observability analysis for continuum robots, emphasizing boundary constraints and sensor placement. While the core principle of examining how small perturbations in the shape state affect sensor measurements remains the same, our approach differs in several key ways: we use local weak observability via Lie derivatives, explicitly incorporate sensor noise through a noise-weighted observability matrix, and evaluate observability via two dedicated indices. Moreover, we integrate this analysis into an EKF framework for real-time state estimation, in contrast to the more structurally focused analysis in [58].
@@ V-B1 Inverse condition number
where $\sigma_{\max}$ and $\sigma_{\min}$ are the maximum and minimum singular values of $\mathcal{O}_{\mathbf{wm}}$, respectively. As a matrix nears rank deficiency, its inverse condition number approaches zero; a value closer to 1 indicates stronger observability. Nonetheless, being dimensionless, it may not always discriminate well among different noise levels [59].
@@ V-B2 Noise amplification index
To overcome this limitation, we also use the noise amplification index [59], defined as
@@ VI-A Computation Time
We benchmark wall-clock per-update latency and scaling within a single Python (NumPy/SciPy) runtime on the same single-threaded machine, so methods are compared under identical interpreter/dependency conditions rather than across languages or hardware. The suite includes a CRT shooting solver and a Cosserat-prior GPR estimator adapted from public references [20, 6], alongside PCK forward kinematics and the proposed IMM–PCK estimator. CRT timings are reported solely as a like-for-like runtime reference to inform loop-rate budgeting—not as an accuracy comparison or a cross-language speed ranking.

### v5

@@ I Introduction
Inspired by biological features such as an elephant’s trunk or an octopus’s tentacles, continuum robots are an unconventional class of robots that have inherent large degrees of flexibility, underactuation, and compliance. Such "soft" features allow them to have safe physical interactions with the environments[hirose1993biologically]. With recent advances in instrument miniaturization, continuum robot designs have been widely used in many minimally invasive surgery applications [simaan2018medical, dupont2022continuum], and their advantages are further leveraged for safe manipulation and access to deep anatomy.
@@ I-A Sensing Technique
Shape sensing is pivotal in this context by supplying the observational data critical for robust shape estimation. A diverse array of sensing techniques is deployed to acquire the shape information. External sensing involves equipping robots with sensors such as Electromagnetic (EM) sensors[song2015electromagnetic, anderson2017continuum, lilge2022continuum, borgstadt2015multi, ataka2016real, shi2016shape], vision-based systems[venkiteswaran2019shape, borgstadt2015multi, loo2019non], or Inertial Measurement Units (IMUs)[cheng2022orientation, peng2024tendon] to obtain pose information externally. Alternatively, intrinsic sensing leverages internal sensors to enhance joint-level information, utilizing encoders or passive cable displacement measurements [orekhov2023lie]. Given their infinite degrees of freedom (DoF), the challenge with continuum robots lies in capturing comprehensive pose information along their entire length with a finite number of sensors. While technologies like Fiber Bragg Grating (FBG) sensors[kim2014optimizing, qiao2019estimating, qiao2021force, al2021fbg, sefati2017highly, shi2016shape] and vision-based systems can map out the full robot shape, constraints such as budget and the specific requirements of the application often preclude their ubiquitous use.
@@ I-B Representations in Modeling
Different shape representation methods yield various formulations of the backbone curvature, which directly impact continuum robots’ computed position and orientation. Representation methods can be categorized as distributed and lumped parameterizations for the backbone curvature[rao2021model, rone2013continuum].
Distributed parameterization methods represent the backbone curvature as a continuous function of the arc length, characterized by an infinite number of parameters. These methods assume that the curvature distribution is variable along the length of the robot. A classic example is the Cosserat rod theory (CRT), which employs distributed parameter modeling to account for the underlying mechanics by solving for positional, force, and moment boundary conditions [jones2009three, till2019real, boyer2020dynamics, renda2018discrete, rucker2011statics].
Lumped parameterization uses a finite set of parameters to describe the backbone shape of continuum robots geometrically. This approach primarily includes three methods: the Pseudo-Rigid-Body Model (PRBM), the constant curvature (CC) assumption, and the modal approach[rao2021model]. PRBM approximates the central backbone shape of continuum robots as a series of rigid links connected by torsion springs[huang20193d, su2009pseudorigid, khoshnam2015robotics, roesthuis2016steering]. The CC model assumes that each segment of the continuum robot follows a constant curvature, typically represented as a circular arc, with a single curvature parameter defining the shape. Due to its simplification and computation efficiency, it is deployed quickly for ideal situations such as uniform mass/material distribution, frictionless transmission, and free from external loading. As a widely adopted kinematics-based approach, The CC model is extensively used in applications ranging from surgical procedures [simaan2004dexterous, sears2006steerable, camarillo2008mechanics, webster2010design] and field operations[hannan2003kinematics] to underwater exploration[sitler2022modular] and aerial maneuvers [zhao2022modular]. In contrast, the modal approach[chirikjian1994modal] represents backbone curvature as a linear combination of modal shape functions (e.g., monomials[wang2019geometric], Euler curves [rao2022shape], Chebyshev polynomials[orekhov2023lie], and polynomial curvature kinematics (PCK)[cheng2022orientation, della2019control]). A modal approach offers the advantage of capturing complex, continuous deformations with a relatively small number of parameters, making it computationally efficient while still maintaining high accuracy in shape representation.
In this work we adopt the polynomial-curvature (PCK) modal family as the lumped parameterization; our experiments bring prior PCK theory into hardware validation under limited/sparse sensing, responding to calls for experimental studies and fair comparisons among polynomial-shape methods [della2019control, sadati2023reduced].
@@ I-C Estimation Approach
Various sensing techniques and representation methods drive multiple estimation approaches tailored to the specific scenario. In static configurations, where the continuum robot remains stationary and its state does not change over time, researchers often employ curve fitting or optimization methods based on modal approaches[kim2014optimizing, song2015electromagnetic, rao2022shape, orekhov2023lie], and PRBM[venkiteswaran2019shape]. Considering model uncertainty and measurement noise, Lilge et al. [lilge2022continuum] and Ferguson et al. [ferguson2024unified] implemented Gaussian Process Regression (GPR) for static shape estimation under arclength domain using CRT.
Dynamic configurations involve continuous state updates as the continuum robot moves and its state changes over time. In such scenarios, stochastic observer-based estimation methods, particularly filtering approaches, are preferred for their ability to optimize state estimation by balancing the influences of process and measurement models, where noise from both sources significantly impacts accuracy and robustness. Although the Extended Kalman Filter (EKF) is a general estimation technique that can be paired with various forward models, most existing continuum robot shape-estimation studies have employed EKF in conjunction with CC or piecewise constant curvature (PCC) models[ataka2016real, loo2019non, peng2024tendon]. In addition to the EKF, alternative filtering methods—such as the Unscented Kalman Filter (UKF) and particle filters—have also been explored [chen2019model, borgstadt2015multi]. Each technique presents trade-offs in computational complexity, robustness, and ease of implementation. For instance, UKF’s sigma-point method can improve estimation in highly nonlinear systems, whereas particle filters can capture non-Gaussian noise at the expense of increased computation time. In certain cases, the EKF remains a better choice when the system model is sufficiently accurate, or when computational resources are limited. Another well-established multi-model estimation approach is the Interacting Multiple Model (IMM) approach [mazor1998interacting, dingler2022state]. Initially developed for tracking targets with Markovian switching dynamics, IMM simultaneously executes multiple Kalman-type filters (e.g., EKF or UKF), each associated with distinct motion models. Estimates from these parallel filters are integrated based on model probabilities updated via measurement innovations, commonly determined through likelihood or Normalized Innovation Squared (NIS) tests. IMM-based approaches are advantageous when a system switches between distinct kinematic regimes or experiences abrupt behavioral changes, conditions frequently encountered in continuum robotics due to varying loads, external interactions, or segmental bending transitions. Despite extensive IMM applications in mobile robotics and trajectory tracking, its deployment in dynamic shape estimation for continuum robots remains scarce, highlighting a valuable research direction in robust multi-model estimation.
@@ III-A Polynomial Curvature Kinematics
A precise representation and model of the shape for continuum robots([della2019control],[cheng2022orientation]) is rooted in PCK. Leveraging this model, all the poses spanning the arclength can be ascertained by integrating a curvature function. The general curvature function is described as an infinite expansion of monomials:
@@ III-A2 Forward kinematics of configuration-to-task space
In this paper, we follow the frame notation of [wang2019geometric] to derive the kinematics (illustrated in Fig. 2), and we list the definitions of all frames in Table I. Because we assume no torsional twist along the backbone, the robot’s shape can be fully described in a single plane (see the right side of Fig. 2). Accordingly, only one bending angle $\delta(t)$ is required to capture the in-plane motion, with Frame $\{1\}$ having its axes $\hat{\mathbf{x}}_{1}$ and $\hat{\mathbf{z}}_{1}$ aligned with the bending plane. Unless stated otherwise, all vectors and parameters are expressed in Frame $\{1\}$.
@@ III-B Shape State Representation
However, this model assumes uniform curvature along the robot’s length and simplifies the representation, which often does not accurately reflect the complex shape dynamics in environments with disturbances. In many constrained or quasi-planar workspaces, such as bench-top manipulation tasks, continuum arms operating within planar cavities, or tendon-driven instruments designed with torsionally stiff backbones, deformation primarily occurs in a single plane. In these contexts, accurate in-plane shape sensing remains essential for navigation and control, while small out-of-plane deflections can be treated as process disturbances within the stochastic framework. Accordingly, this study focuses on planar bending to establish and validate the core observer formulation before extending it to spatial cases in future work. Contemporary systems and surveys report continued use of CC/PCC pipelines in applied settings [gunderman2023non, qiu2025actuator, rogatinsky2023multifunctional, russo2023continuum], while hardware strategies that raise torsional stiffness (thereby suppressing twist in normal use) further support effectively planar operation [santoso2021origami, sun2023enhancing]. Moreover, in multi-segment tendon-driven continuum robots the bending plane angle $\delta$ can be directly computed from cable displacements using established kinematic models [camarillo2008mechanics] and thus does not need to be estimated as an additional state variable. In this work, since we focus exclusively on in-plane bending motions, we omit the bending plane angle $\delta$ from our analysis, thereby simplifying the model to better suit our specific research needs. We intentionally exclude the plane angle $\delta$ from our shape state to maintain a clear focus on how polynomial curvature coefficients map directly to sensor measurements. Incorporating $\delta$ requires augmenting every matrix and Jacobian with rotation operators, significantly increasing the notational overhead. This extra layer, though standard, can obscure the core novelty of using non-constant polynomial curvature within a stochastic observer. By omitting $\delta$, we simplify our presentation of the observer’s theoretical development while still enabling 3D motion in real applications by commanding $\delta$ externally, if desired.
@@ III-C State Transition Model
The scope of our work is confined to evaluating our observer-based shape estimation framework. Consequently, we do not incorporate advanced control strategies but instead opt for a straightforward open-loop control. For simplicity—and without loss of generality—we designate the control input $\mathbf{u}$ as the change in the bending angle at the tip, denoted by $\delta\theta_{e}$. This choice is supported by previous studies, such as Camarillo et al.[camarillo2008mechanics], which demonstrate that in tendon-driven continuum manipulators, the tip bending angle is closely linked to the actuator commands. Similarly, Peng et al.[peng2025dexterous] successfully used the bending angle as a control parameter in an aerial continuum manipulator, illustrating that this input is not only practical but also adaptable across different application domains. Moreover, our IMM-EKF framework is inherently flexible: once a reliable mapping between the chosen control input and the shape state (modal coefficients) is established, the estimation process can accommodate alternative control inputs if desired.
@@ IV-B1 Mitigating Bias in State Dimensionality
In the mixing stage of the IMM, the mixed initial state for each model is derived based on Equation (36), which requires uniform dimensionality across all state vectors and their corresponding covariance matrices. The conventional IMM approach often addresses dimensional discrepancies by setting augmented components in state vectors with smaller dimensions to zero. This leads to potential biases in the mixed initial states of models with larger dimensions. Inspired by [yuan2012multiple], which provides an unbiased mixing solution for two models, we propose an unbiased mixing strategy tailored for multiple models with varying dimensions, illustrated using the case of $\mathbb{M}=\left\{M^{\mathit{(0)}},M^{\mathit{(1)}},M^{(2)}\right\}$.
@@ IV-B2 Designing Adaptive Transition Strategies
The conventional IMM method often struggles to meet performance expectations in dynamic environments due to its reliance on a fixed TPM predicated on prior knowledge. This limitation has spurred researchers to develop methods that enable the TPM to adapt more effectively, not just based on historical data but also incorporating real-time information. We build upon these efforts in our work, especially drawing on the correction functions proposed in [xie2019adaptive] and [lee2023improved]. We leverage these concepts to refine and implement a TPM adjustment mechanism that responds dynamically to current system states, enhancing the IMM’s responsiveness and accuracy in diverse operation scenarios.
@@
We utilize an activating correction function described in [lee2023improved] to reduce lag errors and improve the responsiveness of IMM when the model jumps. The activating function is designed as
@@
Unlike the strategy of using $g_{k}^{(i)}$ at all time steps in [lee2023improved], we introduce the minimum likelihood ratio between models as a prior judging condition to know if the current model jumps. The reason is that this correction function may cause instability in cases with large noises and no dominant model. The likelihood ratio is expressed as
@@ V Observability Analysis
For each model in our IMM framework, an EKF is used to estimate the shape state $\mathbf{m}$. Convergence of the EKF requires that the system be observable. In linear time-invariant (LTI) systems, observability is often checked via the rank of the observability matrix. In nonlinear systems, however, we adopt local weak observability [hermann1977nonlinear], which relies on Lie derivatives to capture how variations in the shape state map to sensor outputs.
Mahoney et al. [mahoney2016inseparable] also employ a rank-based observability analysis for continuum robots, emphasizing boundary constraints and sensor placement. While the core principle of examining how small perturbations in the shape state affect sensor measurements remains the same, our approach differs in several key ways: we use local weak observability via Lie derivatives, explicitly incorporate sensor noise through a noise-weighted observability matrix, and evaluate observability via two dedicated indices. Moreover, we integrate this analysis into an EKF framework for real-time state estimation, in contrast to the more structurally focused analysis in [mahoney2016inseparable].
@@ V-B1 Inverse condition number
where $\sigma_{\max}$ and $\sigma_{\min}$ are the maximum and minimum singular values of $\mathcal{O}_{\mathbf{wm}}$, respectively. As a matrix nears rank deficiency, its inverse condition number approaches zero; a value closer to 1 indicates stronger observability. Nonetheless, being dimensionless, it may not always discriminate well among different noise levels [nahvi1996noise].
@@ V-B2 Noise amplification index
To overcome this limitation, we also use the noise amplification index [nahvi1996noise], defined as
@@ VI-A Computation Time
We benchmark wall-clock per-update latency and scaling within a single Python (NumPy/SciPy) runtime on the same single-threaded machine, so methods are compared under identical interpreter/dependency conditions rather than across languages or hardware. The suite includes a CRT shooting solver and a Cosserat-prior GPR estimator adapted from public references [rao2021model, lilge2022continuum], alongside PCK forward kinematics and the proposed IMM–PCK estimator. CRT timings are reported solely as a like-for-like runtime reference to inform loop-rate budgeting—not as an accuracy comparison or a cross-language speed ranking.